OpenStudy (anonymous):

I have one more question to whomever would love to explain to me how u solve this equation. m=7/8;(5,-7)

OpenStudy (anonymous):

y-(-7)=m(x-5) so you just sub in the gradient and solve

OpenStudy (anonymous):

i know how to write out the equation but after that im lost =(

OpenStudy (anonymous):

What's the question? just to find the equation?

OpenStudy (anonymous):

im sorry its for an online course

OpenStudy (anonymous):

the question is find the equation of the line having the given slope and containing the given point

OpenStudy (anonymous):

You sub it into the point-gradient formula, which is y-y1=m(x-x1) so using the given, sub in your gradient and your y and x-values to get: y-(-7)=(7/8)(x-5) 8(y+7)=7(x-5) 7x-8y-35-56=0 7x-8y-91=0

OpenStudy (anonymous):

Does that help? :I

OpenStudy (anonymous):

im trying to go over it

OpenStudy (anonymous):

You need to find the slope of the original line because the new line is parallel to it and therefore must have the same slope. Then you can write the equation for the new line using the point slope formula \[y-y_0 = m(x-x_0)\] Where y0 and x0 are the y and x coordinates of a point on the line, and m is the slope of the line.

OpenStudy (anonymous):

where do u get 56 from

OpenStudy (anonymous):

8(y+7)=7(x-5) 8y+56=7x-35 0=7x-8y-35-56

OpenStudy (anonymous):

its not the answer

OpenStudy (anonymous):

I just threw everything over to one side..

OpenStudy (anonymous):

answer is 7/8x-

OpenStudy (anonymous):

83/8

OpenStudy (anonymous):

Ohhh you get that answer if you don't get rid of the fraction

OpenStudy (anonymous):

and leave 8 as the denominator

OpenStudy (anonymous):

right whichs is what i wasnt understandin

OpenStudy (anonymous):

When you leave it as a fraction: y-(-7)=(7/8)(x-5) y+7=(7x/8) - 35/8)

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

can someone type out the working, leaving it in fractional form, it's taking so long for me, ha

OpenStudy (anonymous):

y = (7x/8) - (91/7) confirmed

OpenStudy (anonymous):

\[y+7=[7/8](x-5) \rightarrow y = (7x)/8 - (7*5)/8 -7 \] \[\rightarrow y = (7/8)x - 35/8 - 56/8 = (7/8)x - 91/8\]